# IBNorm: Information-Bottleneck Inspired Normalization for Representation Learning

### v1

@@ 5.2 Ablation Study
Table 4: Ablation study of Llama 60M with normalization using different $\lambda$ evaluated on LLM Leaderboard II. We defer the results of each task on Leadboard I to Tab. 9 of Appendix C.1.
Leaderboard II Leaderboard I
Normalization $\lambda$ IFEval BBH MATH GPQA MUSR MMLU-PRO AVG AVG
IBNorm-S 0.5 0.1687 0.2894 0.0008 0.2592 0.3519 0.1162 0.1977 0.2898
4 0.1781 0.2977 0.0000 0.2501 0.3571 0.1091 0.1987 0.2952
8 0.1623 0.2947 0.0008 0.2606 0.3515 0.1185 0.1981 0.2918
IBNorm-L 0.5 0.1577 0.2968 0.0000 0.2710 0.3690 0.1162 0.2018 0.2863
4 0.1899 0.3001 0.0000 0.2772 0.3519 0.1137 0.2055 0.2955
8 0.1364 0.2888 0.0000 0.2450 0.3452 0.1132 0.1881 0.2921
IBNorm-T 0.5 0.1761 0.3024 0.0000 0.2492 0.3469 0.1158 0.1984 0.2881
4 0.1754 0.3013 0.0008 0.2668 0.3598 0.1110 0.2025 0.2937
8 0.1753 0.2933 0.0000 0.2559 0.3631 0.1156 0.2005 0.2898
Table 5: Ablation study of Llama 60M evaluated on LLM Leaderboard II. Due to space limitation, we defer the results of each task on Leadboard I to Tab. 8 of Appendix C.1.
Leaderboard II Leaderboard I
Normalization IFEval BBH MATH GPQA MUSR MMLU-PRO AVG AVG
IBNorm-S* 0.1626 0.2646 0.0000 0.2876 0.3609 0.1114 0.1978 0.2952
IBNorm-S** 0.1376 0.2946 0.0000 0.2776 0.3460 0.1114 0.1945 0.2881
IBNorm-T* 0.1372 0.2880 0.0000 0.2801 0.3519 0.1166 0.1956 0.2936
IBNorm-T** 0.1272 0.2980 0.0000 0.2701 0.3519 0.1166 0.1940 0.2831
IBNorm-L* 0.1374 0.2941 0.0000 0.2793 0.3505 0.1190 0.1967 0.2905
IBNorm-L** 0.1374 0.2841 0.0000 0.2693 0.3505 0.1190 0.1934 0.2859
Effects of compression hyper-parameter $\lambda$. We analyze the effect of the hyperparameter $\lambda$ in our proposed IBNorm, which controls the strength of the compression on intermediate activations. Specifically, we consider $\lambda=0.5,4,8$, corresponding to mild, moderate, and strong compression, respectively. As shown in Tab. 4, the results show that moderate compression ($\lambda=4$) achieves the best performance, striking a balance between preserving task-relevant information and suppressing irrelevant variability. For instance, IBNorm-S achieves the best scores of 0.2952 and 0.1987. We find that overly mild and strong compression lead to suboptimal performance, excessively reducing the representation diversity.
Effects of operation order and affine reparameterization. To better understand the design choices in our proposed IBNorm, we conduct ablation experiments on two critical components: the operation order and the affine reparameterization. As shown in Tab. 5, IBNorm with notation * denotes applying standardization prior to the compression transformation, namely, $\text{IBNorm}(\bm{x};\lambda)\equiv\eta\circ s_{\lambda}\circ\psi\circ\zeta$. This modification leads to only minor fluctuations in performance compared to the baseline: for example, the baseline achieves average scores of 0.2950 and 0.1997, while IBNorm-S* obtains 0.2952 and 0.1978 on Leaderboard I and II, respectively. This shows that we should do compression and then apply standardization for better performance. In addition, IBNorm with notation ** removes the linear affine component, which consistently results in a pronounced degradation across benchmarks. For instance, IBNorm-L achieves 0.2955 and 0.2055, whereas IBNorm-L** drops to 0.2859 and 0.1934 on Leaderboard I and II, showing the importance of affine reparameterization $(\eta).$
@@ 6 Conclusion
We revisited normalization through the lens of IB principle and proposed a principled framework for designing normalization layers. Then, we introduced IBNorm, a family of nonlinear transforms that compress activations toward their mean, yielding more information-rich representations. Theoretically, we showed that IBNorm achieves stronger information-theoretic optimality—attaining higher IB values and tighter generalization bounds—than variance-centric methods. Empirically, IBNorm consistently outperforms BN, LN, and RMSNorm across both vision and language domains.
Limitations. Our experiments are limited to medium-scale LLMs due to computational constraints. Extending evaluations to larger foundation models is an important direction for future work.
Declaration of LLM usage. LLMs were used solely for polishing and formatting text. All research ideas, methods, and experimental results presented are original contributions of the authors.
Ethics statement
The large language models employed in this research may reflect biases or generate sensitive or potentially offensive content, intended solely for academic and scientific purposes. The opinions expressed within generated outputs do not represent the views of the authors. We remain committed to fostering the development of AI technologies which align with ethical standards and reflect societal values.
Reproducibility Statement
We detail our work in the Methodology section (Sec. 4) and describe implementation details in Sec. 5 and Appendix F.
@@ Appendix B NormalNorm
We discuss the difference of our proposed IBNorm and NormalNorm here. We first provide the specific algorithm of NormalNorm in Algorithm 2.
@@
Algorithm 2 Normal Normalization
Input: $\bm{u}=\{u_{i}\}_{i=1}^{N}$
Input: $\bm{y}=\{y_{i}\}_{i=1}^{N}$
Learnable Parameters: $\gamma$, $\beta$
Noise Factor: $\xi\geq 0$
Standardization:
$\hat{u}\leftarrow\frac{1}{N}\sum_{i=1}^{N}u_{i}$
$\hat{\sigma}^{2}\leftarrow\frac{1}{N}\sum_{i=1}^{N}(u_{i}-\hat{u})^{2}$
$c_{i}\leftarrow\frac{u_{i}-\hat{\mu}}{\sqrt{\hat{\sigma}^{2}+\epsilon}}$
Power Transform and Scaled Additive Noise:
@@
with gradient tracking disabled:
$\bar{x}=\frac{1}{N}\sum^{N}_{i=1}x_{i}$
$s=\frac{1}{N}\sum^{N}_{i=1}|x_{i}-\bar{x}|$
sample $z_{i}\sim\mathcal{N}(0,1)$
$v_{i}=x_{i}+z_{i}\cdot s\cdot\xi$
Affine Transform:
$y_{i}\leftarrow\gamma\cdot v_{i}+\beta$
While NormalNorm also attempts to explicitly shape activation distributions, encouraging Gaussian-like features through the power transform with additive noise, this approach has several limitations. In practice, we find that dynamically varying $\lambda$ not only increases computational cost, but can also lead to training instability due to numerical issues, especially in the setting of lower precision like bfloat16. Furthermore, because the power transform modifies both the first- and second-order statistics of the hidden activations, the resulting outputs after the NOP no longer have zero mean and unit variance. As a result, the network cannot fully benefit from traditional normalization techniques, such as improved training stability and scale invariance.
@@ C.1 Results on LLM Leaderboard I
Model Normalization ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K AVG
@@
Model Normalization ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K AVG
@@
Normalization ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K AVG
@@
Normalization $\lambda$ ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K AVG
@@
C.2 Results on ImageNet

### v2

@@ 5.2 Ablation Study
Effects of compression hyper-parameter $\lambda$. We analyze the effect of the hyperparameter $\lambda$ in our proposed IBNorm, which controls the strength of the compression on intermediate activations. We consider $\lambda=0.5,4,8$, corresponding to mild, moderate, and strong compression, respectively. As shown in Tab. 4, the results show that moderate compression ($\lambda=4$) achieves the best performance, striking a balance between preserving task-relevant information and suppressing irrelevant variability. For instance, IBNorm-S achieves the best scores of 0.2952 and 0.1987. We find that overly mild and strong compression lead to suboptimal performance, excessively reducing the representation diversity.
Effects of compression operation and linear affine reparameterization. To further clarify the design choices underlying our proposed IBNorm, we perform ablation studies on two key components: the compression operation and the standardization step. As reported in Tab. 5, the variant of IBNorm marked with an asterisk (*) applies standardization before the compression operation, namely, $\text{IBNorm}(\bm{x};\lambda)\equiv\eta\circ s_{\lambda}\circ\psi\circ\zeta$. This modification leads to only minor fluctuations in performance compared to the baseline: for example, the baseline achieves average scores of 0.2950 and 0.1997, while IBNorm-S* obtains 0.2952 and 0.1978 on Leaderboard I and II, respectively. These results indicate that applying compression operation prior to standardization is preferable for optimal performance. In addition, the variant denoted by a double asterisk (**) removes the linear affine component, which results in a substantial performance drop across benchmarks. For instance, IBNorm-L achieves 0.2955 and 0.2055, whereas IBNorm-L** drops to 0.2859 and 0.1934 on Leaderboard I and II. This pronounced degradation highlights the critical role of the affine reparameterization $(\eta)$ in the IBNorm design.
@@ 6 Conclusion
We revisited normalization through the lens of IB principle and proposed a principled framework for designing normalization layers. We introduced IBNorm, a family of nonlinear transforms that compress activations toward their mean, yielding more information-rich representations. Without relying on explicit IB objectives during training, IBNorm remains model-agnostic and incurs no heavy computational overhead. Theoretically, we showed that IBNorm achieves stronger information-theoretic optimality—attaining higher IB values and tighter generalization bounds—than variance-centric methods. Empirically, IBNorm outperforms BN, LN, and RMSNorm across vision and language domains.
Limitations. Our experiments are limited to medium-scale LLMs due to computational constraints. Extending the evaluation to larger foundation models will be addressed when sufficient computational resources are available.
Impact Statement
The pretrained LLMs employed in this research may reflect biases or generate sensitive or potentially offensive content, intended solely for academic and scientific purposes. The opinions expressed within generated outputs do not represent the views of the authors. We remain committed to fostering the development of AI technologies which align with ethical standards and reflect societal values.
@@ Appendix A IBNorm
A.1 Comparison with Task-Irrelevant Information Compression Methods in Representation Learning
Prior works (Tishby et al., 2000; Tishby and Zaslavsky, 2015; Saxe et al., 2019; Hua et al., 2021; Weng et al., 2023; Hu et al., 2025) explore the IB theory and representation learning methods, such as whitening, and empirically shows how hidden layers with activation nonlinearities and whitening modules aiming to decorrelate features and stabilize training can compress task-irrelevant information.
While these findings provide valuable empirical insights, there are two notable gaps unaddressed: (1) they do not examine how normalization layers influence the decomposition of task-relevant and task-irrelevant mutual information within neural representations, and (2) they do not provide a theoretical framework for how to design the normalization following the IB framework to facilitate more informative and robust representations. Importantly, activation and normalization are conceptually and structurally distinct: In modern LLMs (e.g., LLaMA, GPT), nonlinear activations occur only in the gated branch of the MLP, and do not influence the representations within each transformer block. In these architectures, inserting tanh as an ”activation” would not propagate compression throughout the model, since the gating branch represents only one component of the block.
In contrast, IBNorm applies compression directly the normalization layer, which is invoked throughout the model—after each attention block and MLP block in LLMs. This placement ensures that compression systematically modulates the IB value of the representation at every layer, producing effects that cannot be achieved by modifying the activation function alone. In addition, recent studies on normalization module in Transformer (Zhu et al., 2025; Chen et al., 2025) replace variance-centric normalization layers with point-wise functions such as Dynamic Tanh, showing that the resulting performance gains arise primarily from improved generalization rather than increased fitting capacity. This observation is consistent with the IB principle underlying our work. Our proposed IBNorm is grounded on the IB-theoretic analysis of normalization and demonstrates how different IB-motivated compression operations shape higher-order statistics, aligning normalization behavior with the IB principle.
A.2 Comparison with Training Pipelines Based on Explicit Information Bottleneck Objectives
Methods (Kolchinsky et al., 2019; Alemi et al., 2016; Belghazi et al., 2018) focus on estimating mutual information by parameterizing the IB principle with pretrained neural MI estimators and then guiding the training of deep learning models using the explicit estimated IB objective. These approaches typically rely on encoder-decoder architectures (e.g., VAEs), require large training datasets to train accurate and lightweight MI estimators before training the deep learning models, and introduce additional estimated MI-regularization loss functions during training the deep learning models. These components introduce extra hyperparameters, increase computational cost, and may lead to unstable optimization. Specifically, these approaches face two key limitations when applied to deep learning models, such as autoregressive LLMs where the sequential decoding process complicates the design and increases the inference cost of MI estimators, in practice:
MI Neural Estimator Architecture and Computational Overhead: These methods require additional MI neural estimators with architectural constraints alongside the main network, significantly increasing heavy inference computational cost and complicating optimization. In addition, designing such neural estimators is particularly difficult for autoregressive LLMs due to their sequential decoding characteristics.
Sample Complexity: Accurate MI estimation, as noted in MINE (Belghazi et al., 2018), requires large sample sizes, which is challenging for models with long-context inputs and requires heavy training cost.
IBNorm fills this gap by providing the first theoretical analysis of normalization under the IB principle. Specifically, IBNorm incorporates the IB principle within the normalization operation. It requires no auxiliary MI estimators, no additional IB-based training losses, and no architectural constraints. The compression operation introduced in IBNorm is not an external module but a theoretically grounded reformulation of normalization that increases the IB value of intermediate representations while adding minimal computational overhead.
We highlight the key advantages of IBNorm here:
Systematic Theoretical Grounding: We provide a theoretical analysis of widely used variance-centric normalization methods. This analysis motivates a principled mechanism for designing normalization layers that follow IB principle, as shown in Sec. 4.3.
Computational Efficiency: Unlike (Kolchinsky et al., 2019; Alemi et al., 2016; Belghazi et al., 2018), IBNorm does not rely on computationally heavy MI neural estimators or explicit IB regularizers, making it more efficient and stable during training. We include the comparison of training time and VRAM across IBNorm and baselines in Appendix C.8.
Compatibility: Probabilistic coding method like Structured Probabilistic Coding (SPC) (Hu et al., 2024) introduces an IB-inspired coding mechanism, but it is limited to encoder-only models and does not examine normalization layers. IBNorm is model-agnostic and can be directly integrated into Transformers (e.g., LLaMA, GPT-2) and CNNs (e.g., ResNet) without modifying the training pipeline, as demonstrated in Sec. 5.
A.3 IBNorm Algorithm
@@ Appendix B NormalNorm
We first provide the specific algorithm of NormalNorm in Algorithm 2.
@@
While NormalNorm also attempts to explicitly shape activation distributions, encouraging Gaussian-like features through the power transform with additive noise, this approach has several limitations: 1) In practice, we find that dynamically varying $\lambda$ not only increases computational cost (see Appendix C.8), but can also lead to training instability due to numerical issues, especially in the setting of lower precision like bfloat16. 2) Furthermore, because the power transform modifies both the first- and second-order statistics of the hidden activations, the resulting outputs after the NOP no longer have zero mean and unit variance. As a result, the network cannot fully benefit from traditional normalization techniques, such as improved training stability and scale invariance.
Algorithm 2 Normal Normalization (Eftekhari and Papyan, 2025)
1:Input: $\bm{u}=\{u_{i}\}_{i=1}^{N}$
2:Input: $\bm{y}=\{y_{i}\}_{i=1}^{N}$
3:Learnable Parameters: $\gamma$, $\beta$
4:Noise Factor: $\xi\geq 0$ Standardization:
@@
[…]
9: with gradient tracking disabled:
[…]
12: sample $z_{i}\sim\mathcal{N}(0,1)$
[…]
@@ C.1 Results on LLM Leaderboard I
Model Normalization ARC ($\uparrow$) HellaSwag ($\uparrow$) MMLU ($\uparrow$) TruthfulQA ($\uparrow$) Winogrande ($\uparrow$) GSM8K ($\uparrow$) AVG ($\uparrow$)
@@
Model Normalization ARC ($\uparrow$) HellaSwag ($\uparrow$) MMLU ($\uparrow$) TruthfulQA ($\uparrow$) Winogrande ($\uparrow$) GSM8K ($\uparrow$) AVG ($\uparrow$)
@@
C.2 Results with Statistical Significance
Table 8: Results of Llama models evaluated on LLM Leaderboard I and II.
Model Normalization Leaderboard I ($\uparrow$) Leaderboard II ($\uparrow$)
Llama 60M LayerNorm 0.2910 $\pm$ 0.0020 0.1922 $\pm$ 0.0033
RMSNorm 0.2895 $\pm$ 0.0027 0.1927 $\pm$ 0.0038
NormalNorm 0.2823 $\pm$ 0.0029 0.1895 $\pm$ 0.0053
IBNorm-L 0.2955 $\pm$ 0.0022 0.2055 $\pm$ 0.0031
Llama 130M LayerNorm 0.2919 $\pm$ 0.0021 0.1942 $\pm$ 0.0035
RMSNorm 0.2913 $\pm$ 0.0027 0.1940 $\pm$ 0.0039
NormalNorm 0.2918 $\pm$ 0.0031 0.1927 $\pm$ 0.0055
IBNorm-T 0.2973 $\pm$ 0.0024 0.2133 $\pm$ 0.0037
Llama 350M LayerNorm 0.3062 $\pm$ 0.0020 0.2003 $\pm$ 0.0037
RMSNorm 0.3025 $\pm$ 0.0028 0.2010 $\pm$ 0.0041
NormalNorm 0.3035 $\pm$ 0.0033 0.1962 $\pm$ 0.0058
IBNorm-L 0.3101 $\pm$ 0.0021 0.2116 $\pm$ 0.0047
IBNorm-T 0.3089 $\pm$ 0.0022 0.2140 $\pm$ 0.0050
Llama 1B LayerNorm 0.3138 $\pm$ 0.0024 0.2053 $\pm$ 0.0042
RMSNorm 0.3134 $\pm$ 0.0030 0.2019 $\pm$ 0.0048
NormalNorm 0.3051 $\pm$ 0.0049 0.2006 $\pm$ 0.0067
IBNorm-T 0.3199 $\pm$ 0.0027 0.2162 $\pm$ 0.0053
Table 9: Results of GPT-2 models evaluated on LLM Leaderboard I and II.
Model Normalization Leaderboard I ($\uparrow$) Leaderboard II ($\uparrow$)
GPT-2 124M LayerNorm 0.2888 $\pm$ 0.0017 0.1889 $\pm$ 0.0021
RMSNorm 0.2861 $\pm$ 0.0019 0.1883 $\pm$ 0.0025
NormalNorm 0.2923 $\pm$ 0.0027 0.1983 $\pm$ 0.0031
IBNorm-L 0.2969 $\pm$ 0.0022 0.2068 $\pm$ 0.0025
GPT-2 355M LayerNorm 0.2978 $\pm$ 0.0024 0.1920 $\pm$ 0.0037
RMSNorm 0.2949 $\pm$ 0.0031 0.1948 $\pm$ 0.0042
NormalNorm 0.2953 $\pm$ 0.0035 0.1985 $\pm$ 0.0044
IBNorm-L 0.3034 $\pm$ 0.0028 0.2125 $\pm$ 0.0048
Discussion of Failure Cases. Our evaluation reveals distinct performance characteristics across different task types. While IBNorm consistently improves performance on tasks involving factual retrieval and general language modeling, it achieves suboptimal performance on reasoning-intensive benchmarks such as BBH, GPQA, and MMLU-PRO. These benchmarks place substantial demands on the model’s ability to preserve and integrate intermediate representations across multiple inference steps, particularly in deeper layers. Qualitative analysis shows that models trained with IBNorm tend to produce more concise rationales compared to LayerNorm and RMSNorm baselines. In many challenging cases, the model correctly identifies the key premises and relevant entities, but intermediate inference steps are less explicitly represented, which can affect performance on tasks requiring extended deductive chains. To further characterize this behavior, we examine the entropy of hidden representations and output distributions. We find that IBNorm consistently induces lower activation and token-level entropy in deeper layers for reasoning tasks, indicating a more compact representational structure.
From an information-theoretic perspective, this behavior reflects the inherent trade-off between minimality and sufficiency in the information bottleneck objective, regulated by the hyperparameter $\lambda$. Stronger compression favors representations that are highly predictive of the target output while minimizing redundancy, which is advantageous for retrieval-oriented tasks but may be less aligned with reasoning benchmarks that benefit from maintaining high-entropy intermediate latent states. These observations motivate a layer-dependent normalization strategy as a promising direction for future work. Applying IBNorm in early or intermediate layers can encourage informative representations, while retaining variance-based normalization in deeper layers may help preserve the representational entropy and flexibility necessary for multi-step reasoning.
C.3 Ablation Study on Normalization Structure
@@
Normalization ARC ($\uparrow$) HellaSwag ($\uparrow$) MMLU ($\uparrow$) TruthfulQA ($\uparrow$) Winogrande ($\uparrow$) GSM8K ($\uparrow$) AVG ($\uparrow$)
@@
Table 11: Ablation study of Llama 60M evaluated on LLM Leaderboard II.
Leaderboard II Leaderboard I
Normalization IFEval ($\uparrow$) BBH ($\uparrow$) MATH ($\uparrow$) GPQA ($\uparrow$) MUSR ($\uparrow$) MMLU-PRO ($\uparrow$) AVG ($\uparrow$) AVG ($\uparrow$)
IBNorm-S* 0.1626 0.2646 0.0000 0.2876 0.3609 0.1114 0.1978 0.2952
IBNorm-S** 0.1376 0.2946 0.0000 0.2776 0.3460 0.1114 0.1945 0.2881
IBNorm-T* 0.1372 0.2880 0.0000 0.2801 0.3519 0.1166 0.1956 0.2936
IBNorm-T** 0.1272 0.2980 0.0000 0.2701 0.3519 0.1166 0.1940 0.2831
IBNorm-L* 0.1374 0.2941 0.0000 0.2793 0.3505 0.1190 0.1967 0.2905
IBNorm-L** 0.1374 0.2841 0.0000 0.2693 0.3505 0.1190 0.1934 0.2859
C.4 Ablation Study on $\lambda$
@@
Normalization $\lambda$ ARC ($\uparrow$) HellaSwag ($\uparrow$) MMLU ($\uparrow$) TruthfulQA ($\uparrow$) Winogrande ($\uparrow$) GSM8K ($\uparrow$) AVG ($\uparrow$)
@@
Table 13: Ablation study of Llama 60M with normalization using different $\lambda$ evaluated on LLM Leaderboard II.
Leaderboard II Leaderboard I
Normalization $\lambda$ IFEval ($\uparrow$) BBH ($\uparrow$) MATH ($\uparrow$) GPQA ($\uparrow$) MUSR ($\uparrow$) MMLU-PRO ($\uparrow$) AVG ($\uparrow$) AVG ($\uparrow$)
IBNorm-S 0.5 0.1687 0.2894 0.0008 0.2592 0.3519 0.1162 0.1977 0.2898
4 0.1781 0.2977 0.0000 0.2501 0.3571 0.1091 0.1987 0.2952
8 0.1623 0.2947 0.0008 0.2606 0.3515 0.1185 0.1981 0.2918
IBNorm-L 0.5 0.1577 0.2968 0.0000 0.2710 0.3690 0.1162 0.2018 0.2863
4 0.1899 0.3001 0.0000 0.2772 0.3519 0.1137 0.2055 0.2955
8 0.1364 0.2888 0.0000 0.2450 0.3452 0.1132 0.1881 0.2921
IBNorm-T 0.5 0.1761 0.3024 0.0000 0.2492 0.3469 0.1158 0.1984 0.2881
4 0.1754 0.3013 0.0008 0.2668 0.3598 0.1110 0.2025 0.2937
8 0.1753 0.2933 0.0000 0.2559 0.3631 0.1156 0.2005 0.2898
Table 14: Ablation study of GPT-2 124M with normalization using different $\lambda$ evaluated on LLM Leaderboard I.
Normalization $\lambda$ ARC ($\uparrow$) HellaSwag ($\uparrow$) MMLU ($\uparrow$) TruthfulQA ($\uparrow$) Winogrande ($\uparrow$) GSM8K ($\uparrow$) AVG ($\uparrow$)
IBNorm-T 0.5 0.2652 0.2634 0.2512 0.4896 0.5092 0.0000 0.2964
4 0.2819 0.2693 0.2532 0.4972 0.5185 0.0011 0.3035
8 0.2794 0.2660 0.2522 0.4903 0.5037 0.0000 0.2986
(a) $\lambda=0.5$.
(b) $\lambda=4$.
(c) $\lambda=8$.
Figure 3: Evolution of predictive information $\hat{I}(Y;T_{l})$, task-nuisance information $\hat{I}(T_{l-1};T_{l})$, and the token-level IB value during training of Llama-130M on C4, using IBNorm-T across different hyperparameters $\lambda=0.5,4,8$. Results are reported on the test set.
As shown in Figure 3, we find: 1) Under-compression ($\lambda=0.5$) fails to sufficiently suppress task-irrelevant information, leading to suboptimal IB tradeoffs. 2) Over-compression ($\lambda=8$) starts to suppress task-relevant information, harming predictive performance. 3) Moderate compression ($\lambda=4$) achieves the best IB value, effectively removing nuisance factors while preserving task-useful information.
C.5 Extended Results on ImageNet